### v1

# for Unseen Language Generalization

Abstract
Pretrained language models (PLMs) have shown remarkable generalization toward multiple tasks and languages. Nonetheless, the generalization of PLMs towards unseen languages is poor, resulting in significantly worse language performance, or even generating nonsensical responses that are comparable to a random baseline. This limitation has been a long-standing problem of PLMs raising the problem of diversity and equal access to language modeling technology. In this work, we solve this limitation by introducing LinguAlchemy, a regularization technique that incorporates various aspects of languages covering typological, geographical, and phylogenetic constraining the resulting representation of PLMs to better characterize the corresponding linguistics constraints. LinguAlchemy significantly improves the accuracy performance of mBERT and XLM-R on unseen languages by $\sim$18and $\sim$2%, respectively compared to fully fine-tuned models and displaying a high degree of unseen language generalization. We further introduce AlchemyScale and AlchemyTune, extension of LinguAlchemy which adjusts the linguistic regularization weights automatically, alleviating the need for hyperparameter search. LinguAlchemy enables better cross-lingual generalization to unseen languages which is vital for better inclusivity and accessibility of PLMs.
1 Introduction
Significant advancements in language processing technology have been achieved through the development of PLMs, leading to a commendable proficiency in language comprehension and generation Devlin et al. (2019); Liu et al. (2019); Conneau et al. (2020); Sanh et al. (2022); Lewis et al. (2019); Raffel et al. (2023); Li et al. (2021); Cahyawijaya et al. (2021); Wilie et al. (2020). However, there remains a notable deficiency in the ability of these models to generalize effectively to unseen languages, resulting in a considerable performance reduction of PLMs across thousands of unseen languages. To mitigate this problem, efforts to develop efficient language adaptation approaches are underway, focusing on the incorporation of these unseen languages to PLMs Pfeiffer et al. (2021b); Alabi et al. (2022); Ebrahimi et al. (2022); Goyal et al. (2021).
Incorporating new unseen languages has been a longstanding problem in natural language processing (NLP) research, especially given that most of these unseen languages are low-resource and underrepresented making PLMs difficult to adapt to these languages. MAD-X Pfeiffer et al. (2020b) employs a language adapter to learn new unseen languages by incorporating language adapters that mitigate the risk of forgetting pre-trained knowledge which is known as the curse-of-mulitlinguality. Nonetheless, this approach requires training for generalizing to new unseen languages which makes it costly and difficult to scale to thousands of languages. MAD-G Ansell et al. (2021) and Udapter Üstün et al. (2020) further generalize this approach by utilizing a linguistic-driven contextual parameter generator (CPG) module to generate language-specific parameters allowing the models to generalize to other languages with similar linguistic characteristics. Recently, Rathore et al. (2023) introduced ZGUL, which combines representations over multiple language adapters to generate the unseen language representation. Despite the effectiveness, all these approaches largely rely on two assumptions, i.e., 1) strict categorization of languages and 2) knowing the language category of the query apriori. The first assumption disregards the fact that linguistic phenomena such as code-mixing may occur in the query. While the second assumption might cause performance degradation due to the error propagation from the language identification module Adilazuarda et al. (2023).
To overcome those limitations, in this work, we introduce LinguAlchemy. Unlike adapter-based approaches which isolate the capability to understand different languages over multiple language-specific adapters, LinguAlchemy learns a shared representation across multiple languages, allowing the model to leverage shared knowledge between languages. Such behavior is attainable by eliminating the language-specific module from the model and instead utilizing a regularization to learn language-specific knowledge. LinguAlchemy allows the model to perform inference without knowing the language of the query prior. Our evaluation suggests that LinguAlchemy can improve XLM-R generalization in unseen languages while maintaining the performance on the high-resource languages without requiring information on the language of the query apriori.
In summary, our contributions are as follows:
  1. 1.
    We propose LinguAlchemy, a regularization method that improves unseen language performance on language models and aligns them to arbitrary languages.
  2. 2.
    We demonstrate strong performance on unseen languages for models trained with LinguAlchemy.
  3. 3.
    We introduced a dynamic scaling method to scale the classification and auxiliary loss factors used in the fine-tuning stage.
2 Related Works
PLMs with their transformer-based architectures have been demonstrating exceptional capabilities in language comprehension and generation. These models excel in abstract linguistic generalization by capturing complex linguistic patterns and understanding structural positions and thematic roles, which are crucial for interpreting language semantics. Research in this area Ganesh et al. (2021) has provided critical insights that enable these models to process and generate human language effectively. The studies have explored how these models grasp intricate linguistic features, including syntax and semantics, thereby enhancing their performance across a wide range of language tasks .
In parallel, the development of resources like URIEL vector and lang2vec utility Littell et al. (2017) has been instrumental in extending the reach of multilingual NLP, particularly for less-resourced languages. These tools provide vector representations of languages, leveraging typological, geographical, and phylogenetic data, thus offering a structured approach to understanding linguistic diversity. Complementing this, recent research has conducted a comprehensive survey on the utilization of typological information in NLP, highlighting its potential in guiding the development of multilingual NLP technologies Ponti et al. (2019). This survey emphasized the underutilization of typological features in existing databases and the need for integrating data-driven induction of typological knowledge into machine learning algorithms.
Recent advancements in prefix tuning Li and Liang (2021) and subspace learning Zhang et al. (2020) have contributed significantly to improving generalization in PLMs. These methods focus on learning prefix subspaces to stabilize the direct learning of embeddings, addressing the instability issues present in earlier approaches. Two notable methods are the MAD-X Pfeiffer et al. (2020b) and MAD-G Ansell et al. (2021) frameworks, which employ adapter-based techniques for multi-task cross-lingual transfer, highlighting modularity and parameter efficiency. However, it inherits the limitations of the pretrained multilingual models, such as the limited capacity to adapt effectively to low-resource and unseen languages. Furthermore, while the framework facilitates adaptation to specific target languages, it may bias the model towards these languages, potentially impacting its performance on other languages.
However, despite these advancements, PLMs still face significant challenges in generalizing to unseen languages, particularly when adapting to low-resource and unseen languages. These challenges stem from the vast structural and semantic variation across languages Bender ; Jurafsky and Martin (2019), the scarcity of resources Mohammad (2019); Lewis et al. (2020), and the limitations inherent in the models themselves Lin et al. (2017); Vaswani et al. (2017). This situation highlights the complexity of scaling and generalizing these models effectively and underscores the need for more sophisticated approaches in model training and adaptation to ensure broader and more equitable language coverage.
3 Unseen Languages Adaptation with LinguAlchemy
In this section, we provide an overview of how LinguAlchemy can capture linguistic constraint and how is the intuition behind LinguAlchemy. We also discuss in detail how do we align model representations with the linguistic vector.
3.1 Does Multilingual LMs capture Linguistic Constraint?
In this work, we define the linguistic knowledge as a vector gathered from URIEL vector (Littell et al., 2017). We chose three distinct linguistic knowledge from the database, namely ’syntax_knn’, ’syntax_average’ [^1] In this work, we chose the ’knn’ and ’average syntax features. These include consensus values (like averages) and predicted values (such as kNN regressions based on phylogenetic or geographical neighbors), and ’geo’ features. The choice of ’syntax_knn’ and ’syntax_average’ is motivated by the typological nature of syntax. Syntax in languages varies widely; hence, by using aggregate measures like averages and k-nearest neighbors (kNN), we can capture a more general representation of syntactic features across languages. These features include consensus values, like averages, and predicted values, such as kNN regressions based on phylogenetic or geographical neighbors.
We excluded phonological features, as our focus is not on speech but on textual data. Phonology primarily pertains to spoken languages and would not add significant value to our analysis of written language structures. Additionally, we excluded language family features (’family’). These features are typically binary, indicating whether a language belongs to a particular subfamily. While they can be useful for language identification tasks, they tend to be sparse and may not provide the granularity needed for our study. By focusing on syntax and geographical features, our approach aims to encompass a broader and more nuanced understanding of linguistic variations in a multilingual context.
Syntax Feature
These feature vectors denote a typological feature that is adapted from several sources including World Atlas of Language Structures (WALS), Syntactic Structures of World Languages Collins and Kayne (2009), and short prose descriptions on typological features in Ethnologue Lewis (2009). Syntax vectors captures information about the syntactic properties of languages which derived from large-scale typological databases, which document the structural and semantic variation across different languages. These syntax features in the URIEL vector are utilized to represent languages in vector form, allowing for the analysis and comparison of languages based on their syntactic properties.
Geographical Feature
On the other hand, geographical features represent languages in terms of their geographical properties. The inclusion of ’geo’ features aims to capture geographical attributes of languages. Geographic factors can significantly influence language evolution and structure, making them crucial for understanding linguistic variations. This feature expresses geographical location with a fixed number of dimensions that each represents the “great circle” distance—from the language in question to a fixed point on the Earth’s surface. By incorporating geographical information into language vectors, URIEL and lang2vec provide a more comprehensive view of languages, considering not only their structural and semantic properties but also their geographical context.
3.2 Proof of Concept
Linguistic Separability in LMs
To explore whether multilingual language models (MLMs), such as Multilingual BERT (mBERT), capture the linguistic constraints as defined by the URIEL vectors, we align the mBERT language embeddings to the linguistic knowledge vectors. This projection is quantitatively measured to assess the representation of both seen and unseen languages within the model. The fundamental question we address is the extent to which mBERT’s embeddings correspond to the typological and geographical features encapsulated in the URIEL vectors.
Formally, we define the language representation alignment in Algorithm 1, where $F_{U}$ represents the features extracted from URIEL, $S$ is the set of sentence representations, $H_{x}$ and $N_{x}$ are the hidden states and number of attention-masked tokens for a sentence $x$, respectively. The matrix $W$ is used for the linear projection, and $A$ holds the final aligned representations. Algorithm 1 outlines the process for aligning language representations we use in Figure 1. It leverages the URIEL database for linguistic features, processes sentences through a language model ($\Theta$), and aligns these with mBERT representations ($M$). The algorithm iteratively updates transformation parameters ($W$ and $b$) through a training loop to minimize the loss between the projected mBERT representations and the target sentence representations in set $S$, thus achieving aligned language representations ($A$).
Algorithm 1 Language Representation and Alignment Process
0:  Dataset $D$, URIEL database $U$, Language Model $\Theta$, mBERT representations $M$
0:  Aligned Language Representations $A$
  $F_{U}\leftarrow$ ExtractFeatures($U$)
  $S\leftarrow\{\}$
  for each sentence $x$ in $D$ do
     $H_{x}\leftarrow$ GetLastHiddenStates($x$, $\Theta$)
     $N_{x}\leftarrow$ CountAttentionMasked($x$)
     $R_{x}\leftarrow\frac{\textsc{Sum}(H_{x})}{N_{x}}$
     $S\leftarrow S\cup\{R_{x}\}$
  end for
  $W,b\leftarrow$ InitializeParameters()
  for each training epoch do
     $P_{U}\leftarrow(W\times S)+b$
     $loss\leftarrow$ ComputeLoss($P_{U}$, $F_{U}$)
     $W,b\leftarrow$ UpdateParametersWithConstraint($W,b,loss$)
  end for
  $A\leftarrow\{\}$
  for each sentence representation $s$ in $S$ do
     $A_{m}\leftarrow(W\times s)+b$
     $A\leftarrow A\cup\{A_{m}\}$
  end for
Figure 1 presents a visual analysis facilitated by UMAP McInnes et al. (2018), showing the correlation between mBERT language representation and the linguistic vectors from the URIEL database ($R^{2}$ = 0.816). By leveraging UMAP, the plot accentuates the principal variances within the joint feature space of the embeddings and vectors. The spatial representation of languages on this plot mirrors their linguistic and geographical relatedness, as encapsulated by mBERT. This visualization underscores the model’s ability to mirror linguistic typologies, with languages sharing common roots such as ’de-DE’ and ’nl-NL’ naturally clustering together. The density and arrangement of these clusters potentially reflect mBERT capacity to capture and represent language family traits. Conversely, the presence of sparser clusters or outliers prompts a closer examination of mBERT’s coverage and consistency in representing diverse linguistic features.
Figure 1: Alignment between mBERT Representation with URIEL Language Representation. The green-shaded areas indicate the sentence representations of mBERT while the brown dots represent the URIEL representations of the corresponding language.
3.3 [Uncaptioned image] LinguAlchemy
We introduce LinguAlchemy as an approach that intuitively aligns model representations with linguistic knowledge, leveraging URIEL vectors. This is operationalized through an auxiliary loss function, involving the training process with a nuanced understanding of linguistic characteristics.
In LinguAlchemy, we enhance the fine-tuning of encoder models such as mBERT for downstream tasks by not only using the regular classification loss but also introducing a novel linguistic regularization term. This is achieved through the implementation of a URIEL loss, designed to align the model’s representations with linguistic knowledge derived from URIEL vectors. Specifically, this process involves applying a linear projection to the model’s pooled output, which aligns it with the URIEL vector space. The URIEL loss is quantified as the mean squared error (MSE) between the projected model outputs and the corresponding URIEL vectors. This dual approach, combining classification loss and URIEL loss, allows for a more linguistically informed model training, enhancing the model’s ability to capture and reflect complex linguistic patterns and relationships.
$L_{uriel}(R,U)=\frac{1}{N}\sum_{i=1}^{N}\|R_{i}-U_{i}\|^{2}$
where $R$ represents the model-generated representations, $U$ denotes the URIEL vectors, and $N$ is the number of data points.
Note that there may be discrepancies between the scales of the standard classification loss and the URIEL loss. To address this, we introduce an optional hyperparameter, denoted as $\lambda$, to scale the URIEL loss appropriately.
Dynamic Scaling Approaches
In addition to the fixed scaling factor, we also explore dynamic adjustment of this scaling factor at each training step. This aims to maintain a balance between the classification and URIEL losses, and even considers making the scale trainable. The final loss formula when training with LinguAlchemyis given by:
$L=\lambda_{cls}*L_{cls}+\lambda_{uriel}*L_{uriel}(R,U)$
We define two methods to implement dynamic scaling:
  1. 1.
    AlchemyScale: This method dynamically adjusts the scaling factor $\lambda$ during training. It initiates with scaling factors set relative to the mean of initial losses, ensuring proportional importance to each loss component. Subsequently, these factors are updated periodically using an Exponential Moving Average (EMA) method, with a conservative adjustment for specific loss components to maintain stability. This approach ensures an optimal balance between different loss components, adjusting the contribution of each to the total loss based on predefined criteria.
  2. 2.
    AlchemyTune: Here, $\lambda$ is conceptualized as a trainable parameter within the model’s architecture. Initialized as part of the model’s parameters, $\lambda$ undergoes optimization during the training process. This method applies the scaling factors to loss components, and an additional mini_loss—representing the deviation of the sum of scaling factors from unity—is computed. This mini_loss is used for backpropagation, enabling the scaling factors to adapt based on the training dynamics and dataset specifics.
Both methods aim to enhance model performance by dynamically and intelligently scaling loss components, with the first method relying on predefined, periodically updated scaling mechanisms, and the second integrating the scaling factor into the model’s learning parameters for adaptive adjustments.
4 Experiment Setting
Datasets
In our experiments, we utilize the MASSIVE Xu et al. (2022), which is a comprehensive collection of multilingual data incorporating intent classification tasks. This dataset is particularly notable for its inclusion of a diverse range of languages, with various language families, genera, and scripts. Specifically, we split MASSIVE into 25 languages that are ’seen’ during finetuning and the rest 27 languages that are ’unseen’, which we exclusively used for evaluation. This splitting is based on the language adapters availability as outlined in the prior research of (Pfeiffer et al., 2020a), which we utilized in the AdapterFusion experiment for our baseline model. For a detailed breakdown of the languages used, including their respective families, genera, and script can be found in Appendix A.
Models
Our research employs two state-of-the-art language models: Multilingual BERT Base (mBERTBase) and XLM-Roberta Base (XLM-RBase). These models are chosen for their robust performance across a wide range of multilingual NLP tasks, making them ideal for our intent classification and language identification objectives.
Hyperparameters
The following hyperparameters are used in our training process:
  • Learning rate: $5\times 10^{-5}$
  • Training epochs: 30
  • Performance metric: Accuracy
  • Early stopping criterion: 3 epochs without improvement
5 Results and Discussion
mBERT XLM-R
Language AF Full FT Ours Full FT Ours
am-ET* 4.61% 5.45% 58.07% 78.55% 77.00%
cy-GB 25.05% 23.81% 30.01% 64.39% 68.96%
af-ZA 57.67% 52.72% 50.16% 82.72% 75.73%
km-KH* 7.83% 8.34% 59.85% 84.57% 78.70%
sw-KE 22.16% 19.77% 54.93% 58.07% 74.90%
mn-MN* 27.57% 27.17% 57.38% 87.53% 76.27%
tl-PH 40.25% 37.46% 66.47% 85.91% 80.42%
kn-IN 40.99% 34.23% 67.84% 80.46% 81.19%
te-IN 34.40% 35.27% 71.88% 84.57% 82.59%
sq-AL 49.53% 44.75% 70.66% 67.92% 82.17%
ur-PK 47.11% 42.80% 69.43% 73.60% 81.81%
az-AZ 63.82% 61.57% 69.18% 80.16% 82.03%
ml-IN 35.84% 27.71% 67.83% 78.92% 81.80%
ms-MY 65.80% 66.48% 67.85% 82.99% 81.80%
ca-ES 73.07% 73.10% 68.37% 87.36% 82.04%
sl-SL 49.29% 47.18% 68.52% 86.28% 82.15%
sv-SE 64.09% 60.05% 68.36% 85.37% 82.20%
ta-IN 41.66% 34.87% 68.63% 84.43% 82.38%
nl-NL 69.97% 70.65% 68.62% 81.98% 82.34%
it-IT 71.86% 70.78% 68.14% 78.31% 82.14%
he-IL 51.24% 48.18% 68.07% 88.74% 82.25%
pl-PL 62.27% 59.99% 67.05% 61.30% 81.60%
da-DK 71.25% 71.72% 66.39% 76.46% 81.43%
nb-NO 68.76% 68.53% 65.71% 78.18% 81.33%
ro-RO 58.68% 54.20% 64.93% 82.82% 81.34%
th-TH 30.43% 24.21% 64.37% 73.30% 81.05%
fa-IR 59.35% 56.89% 64.38% 77.20% 81.04%
Average 47.95% 45.48% 64.19% 78.97% 80.32%
Table 1: Performance of LinguAlchemy in MASSIVE dataset for unseen languages. These models have not seen these languages during fine-tuning. For languages in *, mBERT model also never seen the languages during pre-training.
5.1 LinguAlchemy Performance
To evaluate the effectiveness of our proposed technique on unseen languages, we trained mBERT and XLM-R on the MASSIVE dataset. Specifically, we train on 25 languages and test on 27 different, unseen languages. Our results are summarized in Table 1. We compared our method with zero-shot generalization, where the model is fully tuned on seen languages and then tested on unseen languages (referred to as Full FT in the Table). Additionally, we explored AdapterFusion Pfeiffer et al. (2021a) as another baseline (AF in the Table). AdapterFusion has shown better adaptation to unseen languages than naive zero-shot generalization. Unfortunately, many language adapters that we need for AdapterFusion is not available for XLM-R.
From Table 1, it is shown that LinguAlchemy achieves better generalization for unseen languages. We observed a significant improvement for mBERT and a modest average improvement for the stronger XLM-R model. For mBERT, LinguAlchemy  can significantly increase performance in truly unseen languages of am-ET, km-KH, mn-MN, in which mBERT has never seen during the pre-training stage nor fine-tuning. These findings show that LinguAlchemy can be useful in truly zero-shot settings.
We note that LinguAlchemy tends to even out the performance across all unseen languages, leading to a massive boost in weaker languages such as cy-GB or sw-KE. However, in languages where zero-shot performance is already strong, LinguAlchemy does not seem to provide benefits and in some cases degrade performance. This is more evident in XLM-R, where LinguAlchemy flattens performance to the 80-82% range, in which many of zero-shot performance is already reached that level or more. Regardless, the potential of our method is clear, showing it to be beneficial in cases where zero-shot performance is poor.
Figure 2: Average performance of unseen languages under dynamic URIEL loss scales
5.2 Effect of Scaling URIEL loss
The classification and URIEL losses are not on the same scale. Therefore, simply adding both losses together means that the model will give more weight to the loss with the higher magnitude. When observing both the classification and URIEL losses during the early stages of training, we note that the classification loss is around 10 times larger than the URIEL loss. In this part, we explore the effect of different scaling factors for the URIEL loss.
Constant Scaling
We explore consistently scaling up the URIEL loss across various scaling factors. The results can be seen in Figure 2. It is important to note that, as we use the scale-invariant optimizer AdamW, we don’t have to worry about gradients becoming too large due to extremely large losses. Generally, we observe that a scaling factor of 10x slightly outperforms other scaling factors, and the performance appears to decline with higher scale factors.
Dynamic and Trainable Scaling
One issue with introducing a scaling factor is the addition of another tunable hyperparameter. Intuitively, we might aim for a balanced weight between the classification and URIEL losses. Therefore, instead of expensively testing different scaling factors, an adaptive scaling factor might be more cost-effective and beneficial.
Here, we explore two ideas: dynamic and trainable factors. The results of these approaches can be seen in Table 2.
URIEL scaling mBERT XLM-R
Constant 10x 64.68% 80.32%
Dynamic 62.97% 80.43%
Trainable 63.24% 79.10%
Table 2: Average performance of unseen languages under dynamic URIEL loss scaling factors
Interestingly, these dynamic scale factors do not significantly outperform a constant factor. In contrast, a 10x scaling achieves the best performance in mBERT, while dynamic scaling barely outperforms the 10x scaling in XLM-R. Therefore, in a limited budget scenario, a suggested 10x scaling factor should suffice, and one may explore different scaling factors given more computational resources.
5.3 Seen Language Performance
Performance mBERT XLM-R
Full FT Ours Full FT Ours
Unseen Lgs. 45.48% 64.68% 78.97% 80.43%
Seen Lgs. 84.52% 67.45% 86.45% 81.05%
Average 64.25% 66.01% 82.56% 80.62%
Table 3: Average performance of LinguAlchemy on unseen vs seen languages
One drawback of LinguAlchemy that we have noticed is the sacrifice in performance for seen languages. This result is related to our findings in unseen languages, where LinguAlchemy flattens out the performance across all languages, thus being extremely beneficial to poorly performing ones. Additionally, we note that the performance of seen languages is flattened to a similar level, which, most of the time, is worse than that of standard in-language fine-tuning. The compiled results can be seen in Table 3
Based on this finding, LinguAlchemy is suitable for enhancing the performance of extremely low-resource languages where standard cross-lingual zero-shot fine-tuning does not improve performance. In cases where training data is available, normal fine-tuning is a better choice. We are exploring why LinguAlchemy does not help with seen languages and how to boost the performance of seen languages as future work. Nevertheless, our method is still beneficial in under-resourced settings where multilingual models perform poorly.
6 Conclusion
We introduced LinguAlchemy, a novel approach that demonstrates strong performance across 27 unseen languages in a 60-class intent classification task. Our method hinges on the integration of linguistic knowledge through the URIEL vectors, enhancing the language model’s ability to generalize across a diverse set of languages. We also proposed AlchemyScale and AlchemyTune, which employs a hyperparameter search for the URIEL scaling factor. This is achieved by two key strategies: (1) weight-averaging classification and URIEL loss, and (2) learning to balance the scale between classification and URIEL loss, thus ensuring a more adaptable and robust model performance.
Limitations
Despite the promising results achieved by LinguAlchemy in enhancing the performance across a wide range of unseen languages in intent classification tasks, there are notable limitations in our approach. Firstly, the performance on seen languages remains suboptimal, suggesting a potential area of improvement, possibly through methods such as weight freezing. Secondly, we observe a trade-off where enhancing generalization capabilities seems to come at the cost of reduced accuracy in seen languages. This warrants further investigation into methodologies that can balance performance across both seen and unseen languages more effectively. Additionally, our research is currently limited to intent classification tasks, and extending this approach to a broader range of NLP tasks could provide deeper insights into its versatility and applicability. Another limitation lies in the selection of URIEL features based on theoretical reasoning; while our choice of syntax, geography, and language family features is grounded in Chapter 3, empirical testing with alternative or additional features could uncover opportunities for further refinement of the model. Addressing these limitations could significantly enhance the generalizability and overall performance of multilingual models in NLP applications.
Appendix A Languages in Dataset
The MASSIVE Dataset or Multilingual Amazon SLU Resource Package (SLUPR) which encompasses slot-filling, intent classification, and Virtual Assistant Evaluation tasks, is a new dataset containing about 1 million realistic data examples shaped from utterances by annotators. This dataset spans 51 languages, 60 intents, 55 slot types, and 18 domains.
Table 4: Statistics and description of the dataset used Xu et al. (2022). The dataset used is a subset of the MASSIVE dataset, selecting 25 different seen languages.
Code Name Script Genus Code Name Script Genus
ar-SA Arabic Arab Semitic is-IS Icelandic Latn Germanic
bn-BD Bengali Beng Indic ka-GE Georgian Geor Kartvelian
el-GR Greek Grek Greek km-KH Khmer Khmr Khmer
en-US English Latn Germanic lv-LV Latvian Latn Baltic
es-ES Spanish Latn Romance ml-IN Malayalam Mlym Southern Dravidian
fa-IR Persian Arab Iranian nb-NO Norwegian Latn Germanic
fr-FR French Latn Romance ro-RO Romanian Latn Romance
he-IL Hebrew Hebr Semitic sl-SI Slovenian Latn Slavic
hu-HU Hungarian Latn Ugric ur-PK Urdu Arab Indic
hy-AM Armenian Armn Armenian zh-CN Mandarin Hans Chinese
id-ID Indonesian Latn Malayo-Sumbawan zh-TW Mandarin Hant Chinese
Code Name Script Genus Code Name Script Genus
af-ZA Afrikaans Latn Germanic my-MM Burmese Mymr Burmese-Lolo
am-ET Amharic Ethi Semitic nl-NL Dutch Latn Germanic
az-AZ Azerbaijani Latn Turkic pl-PL Polish Latn Slavic
cy-GB Welsh Latn Celtic pt-PT Portuguese Latn Romance
da-DK Danish Latn Germanic ru-RU Russian Cyrl Slavic
de-DE German Latn Germanic sq-AL Albanian Latn Albanian
fi-FI Finnish Latn Finnic sv-SE Swedish Latn Germanic
hi-IN Hindi Deva Indic sw-KE Swahili Latn Bantoid
ja-JP Japanese Jpan Japanese ta-IN Tamil Taml Southern Dravidian
kn-IN Kannada Knda Southern Dravidian te-IN Telugu Telu South-Central Dravidian
ko-KR Korean Kore Korean th-TH Thai Thai Kam-Tai
mn-MN Mongolian Cyrl Mongolic vi-VN Vietnamese Latn Viet-Muong
ms-MY Malay Latn Malayo-Sumbawan
Table 4: Statistics and description of the dataset used Xu et al. (2022). The dataset used is a subset of the MASSIVE dataset, selecting 25 different seen languages.
Table 5: Statistics and description of the dataset used Xu et al. (2022). The dataset used is a subset of the MASSIVE dataset, selecting 27 different unseen languages.