### v6

# for Unseen Language Generalization

Abstract
Pretrained language models (PLMs) have become remarkably adept at task and language generalization. Nonetheless, they often fail when faced with unseen languages. In this work, we present LinguAlchemy, a regularization method that incorporates various linguistic information covering typological, geographical, and phylogenetic features to align PLMs representation to the corresponding linguistic information on each language. LinguAlchemy significantly improves the performance of mBERT and XLM-R on low-resource languages in multiple downstream tasks such as intent classification, news classification, and semantic relatedness compared to standard approach and displaying a high degree of unseen language generalization. We further introduce AlchemyScale and AlchemyTune, extension of LinguAlchemy which adjusts the linguistic regularization weights automatically, alleviating the need for hyperparameter search.
1 Introduction
Significant advancements in language processing technology have been achieved through the development of PLMs with their impressive capability in language comprehension and generation Devlin et al. (2019); Liu et al. (2019); Lewis et al. (2019); Li et al. (2021); Sanh et al. (2022); Raffel et al. (2023). The development has been further expanded to non-English languages Conneau et al. (2020); Martin et al. (2020); Wilie et al. (2020); Kakwani et al. (2020); Cahyawijaya et al. (2024). However, there is still a gap in these models’ ability to generalize effectively to low-resource and unseen languages, although there have been a numerous work in the field Pfeiffer et al. (2021b); Goyal et al. (2021); Alabi et al. (2022); Ebrahimi et al. (2022); Yong et al. (2023).
Figure 1: LinguAlchemy enhances performance in unseen languages by allowing the model to predict the linguistic vector and then fitting it via a similarity loss towards the specific language’s URIEL vector.
Previous approaches Rathore et al. (2023); Üstün et al. (2022); Pfeiffer et al. (2020b); Ansell et al. (2021) are based on the assumption that disregards the fact that in a real-world application, there is usually no language information from the user, highlighting the importance of the multilingual robustness of a language model. The second assumption might cause performance degradation due to the error propagation from the language identification module Adilazuarda et al. (2023). However, these methods inherit the limitations of the pretrained multilingual models, such as the limited capacity to adapt effectively to low-resource and unseen languages. Furthermore, while the framework facilitates adaptation to specific target languages, it may bias the model towards them, potentially impacting its performance on other languages.
In this work, we introduce LinguAlchemy, a novel method that incorporates a unified representation across multiple languages to enable the model for utilizing the shared linguistic knowledge. Our approach differs from adapter-based approaches which often segment language understanding into multiple, isolated language-specific modules. Instead, we developed a regularization technique that utilizes linguistic information directly into the model’s architecture, allowing for language-agnostic inference. Our evaluations demonstrate that LinguAlchemy not only enhances generalization capabilities of mBERT Devlin et al. (2018) and XLM-R Conneau et al. (2020) on unseen languages but also upholds robust performance across high-resource languages, all without prior knowledge of the query’s language.
Our strategy aims to refine cross-lingual generalization by leveraging linguistic features encapsulated in URIEL vectors. We hypothesize that languages with similar syntactic and geographical characteristics can benefit from shared representational frameworks, significantly boosting performance in multilingual settings. This approach is particularly beneficial in contexts where language resources are limited.
In summary, our contributions are as follows:
  1. 1.
    We propose LinguAlchemy, a regularization method that utilizes geographical and syntactic information to foster the models’ unified representation.
  2. 2.
    LinguAlchemy does not require any architectural change and can be adapted to different tasks and models.
  3. 3.
    We demonstrate strong performance on 50+ languages across three diverse datasets and tasks (intent classification, news classification, and semantic relatedness) for models trained with LinguAlchemy, including the languages that are not seen during pretraining.
  4. 4.
    We introduce two automatic hyperparameter search methods to scale the classification and auxiliary loss factors used in the fine-tuning stage, namely dynamiclearn and dynamicscale.
2 Related Work
PLMs with their transformer-based architectures have been demonstrating exceptional capabilities in language comprehension and generation Ganesh et al. (2021). Rathore et al. (2023) have explored how these models learn intricate linguistic features, including syntax and semantics to enhance their performance across a wide range of language tasks.
Incorporating new unseen languages has been a longstanding problem in the multilingual research, MAD-X Pfeiffer et al. (2020b) employ a language adapter to learn new unseen languages using language adapters that mitigate the risk of forgetting pre-trained knowledge, which is known as the curse-of-multilinguality Conneau et al. (2020). Nonetheless, this approach requires training for generalizing to new unseen languages, which makes it costly and difficult to scale to thousands of languages. MAD-G Ansell et al. (2021) and Udapter Üstün et al. (2020) further generalize this approach by utilizing a linguistic-driven contextual parameter generator (CPG) module to generate language-specific parameters, allowing the models to generalize to other languages with similar linguistic characteristics. Recently, Rathore et al. (2023) introduce ZGUL, which combines representations over multiple language adapters and adding linguistic vector information to generate the unseen language representation. Despite the effectiveness, all these approaches rely on two assumptions: (1) strict categorization of languages and (2) knowing the language category of the query apriori—our definition of “a priori categorization" as incorporating language-specific information into the model.
In parallel, the development of linguistically-driven resources such as the URIEL vector and the lang2vec utility Littell et al. (2017) has been notable in extending multilingual NLP research, particularly for less-resourced languages, by providing methods to represent and compare the structured lingustic features across different languages. Complementing this, Ponti et al. (2019) pointed out the underexplored typological features in existing approaches and the need for integrating data-driven methods of typological knowledge into language models. Previous studies have focused on extending multilingual NLP using URIEL vectors Lauscher et al. (2020); Lin et al. (2019); Tan et al. (2019); Oncevay et al. (2020), but none of their approaches use URIEL vectors for alignment during the finetuning process of language models.
However, PLMs still face significant challenges in generalizing to unseen languages, particularly when adapting to low-resource and unseen languages. These challenges stem from the vast structural and semantic variation across languages Bender (2011); Jurafsky and Martin (2019), the scarcity of resources Mohammad (2019); Lewis et al. (2020), and the limitations inherent in the models themselves Lin et al. (2017). This situation highlights the complexity of generalizing these models effectively to a broader and scope of language. This situation highlights the complexity of generalizing these models effectively to a broader scope of languages, making it difficult for them to perform well on languages they haven’t been specifically trained on.
3 Unseen Languages Adaptation with LinguAlchemy
In this section, we provide an overview of how LinguAlchemy captures linguistic constraints and explain the intuition behind it. We also discuss in detail how we align model representations with the linguistic vectors.
3.1 Does Multilingual LMs capture Linguistic Constraints?
We define the linguistic knowledge as a vector gathered from URIEL vector (Littell et al., 2017). We chose three distinct linguistic knowledge from the database, namely ’syntax_knn’, ’syntax_average’ [^1] In this work, we chose the ’knn’ and ’average syntax features. These include consensus values (like averages) and predicted values (such as kNN regressions based on phylogenetic or geographical neighbors), and ’geo’ features. The choice of ’syntax_knn’ and ’syntax_average’ is motivated by the typological nature of syntax. Syntax in languages varies widely; hence, by using aggregate measures like averages and k-nearest neighbors (kNN), we can capture a more general representation of syntactic features across languages. Note that in our experiments, we excluded phonological features and language family attributes from our analysis as they are less relevant to textual data and have limited granularity for understanding linguistic variations.
Syntax Features
These feature vectors denote a typological feature that is adapted from several sources including World Atlas of Language Structures (WALS) Dryer and Haspelmath (2013), Syntactic Structures of World Languages Collins (2010), and short prose descriptions on typological features in Ethnologue Lewis (2009).Syntax vectors capture information about the syntactic properties of languages, derived from large-scale typological databases that document the structural and semantic variation across different languages. These syntax features in URIEL are utilized to represent languages in vector form that allows the analysis and comparison of languages based on their syntactic properties.
Figure 2: Alignment between mBERT Representation with URIEL Language Representation. The green-shaded areas indicate the sentence representations of mBERT while the brown dots represent the URIEL representations of the corresponding language.
Geographical Features
On the other hand, geographical features represent languages in terms of their geographical properties. The inclusion of ‘‘geo" features aims to capture geographical attributes of languages. This feature expresses geographical location with a fixed number of dimensions that each represents the “great circle" distance—from the language in question to a fixed point on the Earth’s surface. By incorporating geographical information into language vectors, URIEL and lang2vec provide a more comprehensive view of languages, considering not only their structural and semantic properties but also their geographical context.
3.2 Proof of Concept
Linguistic Separability in LMs
We investigate whether PLMs like mBERT Devlin et al. (2018) can capture linguistic constraints by aligning mBERT language embeddings with URIEL vectors to assess how they represent seen and unseen languages. This includes examining how well mBERT’s embeddings correspond to the typological and geographical features detailed in URIEL. In Figure 2, sentence embeddings (green dots) from mBERT, derived from the last hidden state of multilingual training data, and URIEL vectors (brown dots)—structured representations from the URIEL database—are projected into the same space. A matrix $W$ is used to linearly project sentence embeddings, minimizing the mean squared error with URIEL vectors. This alignment is showcased in Figure 2 using UMAP for visualization purpose.
Figure 2 presents a visual analysis facilitated by UMAP McInnes et al. (2018), showing the correlation between mBERT language representation and the linguistic vectors from the URIEL database ($R^{2}$ = 0.816). By leveraging UMAP, the plot highlight the principal variances within the joint feature space of the embeddings and vectors. The spatial representation of languages on this plot mirrors their linguistic and geographical relatedness, as encapsulated by mBERT. This visualization shows the model’s ability to mirror linguistic typologies, with languages sharing common roots such as ’de-DE’ and ’nl-NL’ naturally clustering together. The density and arrangement of these clusters potentially reflect mBERT capacity to capture and represent language family traits. Conversely, the presence of sparser clusters or outliers requires us to carefully check mBERT’s coverage and consistency in representing different linguistic features. We also formally defined the language representation alignment in Algorithm 1.
Model Feature Type Acc.(%)
mBERT geo 66.14
syntax_avg+geo 65.62
syntax_avg 66.41
syntax_knn+geo 66.08
syntax_knn 66.41
syntax_knn+syntax_avg+geo 66.47
syntax_knn+syntax_avg 66.41
XLM-R geo 80.16
syntax_avg+geo 80.54
syntax_avg 80.76
syntax_knn+geo 80.48
syntax_knn 80.32
syntax_knn+syntax_avg+geo 80.80
syntax_knn+syntax_avg 80.78
Table 1: Linguistic vector ablation experiment (the highest accuracy for each model is highlighted in bold).
Table 1 shows the results of the linguistic vector ablation experiment for mBERT and XLM-R, testing different feature combinations. For mBERT, individual features like syntax_avg and syntax_knn perform similarly ( 66.41%), with only minor improvement (66.47%) when combined with geo. In contrast, XLM-R benefits more from feature combinations, achieving its highest accuracy (80.80%) when syntax_knn, syntax_avg, and geo are combined. Even individually, syntax_avg (80.76%) performs well for XLM-R, highlighting the model’s stronger ability to leverage syntactic information. These results suggest that combining syntax and geographical features yields optimal performance, especially for XLM-R, and this combination will be used in subsequent experiments.
3.3 [Uncaptioned image] LinguAlchemy
We introduce LinguAlchemy as an approach that intuitively aligns model representations with linguistic knowledge, leveraging URIEL vectors. This approach is applied through an auxiliary loss function that is involved in the training process with an added information of linguistic characteristics in the form of URIEL vector.
In LinguAlchemy, we enhance the fine-tuning of encoder models such as mBERT for downstream tasks by not only using the regular classification loss but also introducing a novel linguistic regularization term. This is achieved through the implementation of a URIEL loss, designed to align the model’s representations with linguistic knowledge derived from URIEL vectors. Specifically, this process involves applying a linear projection to the model’s pooled output, which aligns it with the URIEL vector space. The URIEL loss is quantified as the mean squared error (MSE) between the projected model outputs and the corresponding URIEL vectors. This dual approach allows for a more linguistically informed model training and increase the model’s ability to capture and reflect complex linguistic patterns.
$\mathcal{L}_{uriel}(Z,U)=\frac{1}{N}\sum_{i=1}^{N}\|Z_{i}-U_{i}\|^{2},$ (1)
LinguAlchemy is represented by equation 1 where $Z$ represents the model-generated representations, $U$ denotes the URIEL vectors, and $N$ is the number of data points. To generate the model representation, we take the output representation from the $CLS$ token and multiply it with a new, trainable projection layer to transform the vector size so that they are compatible.
Formally, we define the language representation alignment in Algorithm 1, where $F_{U}$ represents the features extracted from URIEL, $S$ is the set of sentence representations, $H_{x}$ and $N_{x}$ are the hidden states and number of attention-masked tokens for a sentence $x$, respectively. The matrix $W$ is used for the linear projection, and $A$ holds the final aligned representations. Algorithm 1 outlines the process for aligning language representations we use in Figure 2. It leverages the URIEL database for linguistic features, processes sentences through a language model ($\Theta$), and aligns these with mBERT representations ($M$). The algorithm iteratively updates transformation parameters ($W$ and $b$) through a training loop to minimize the loss between the projected mBERT representations and the target sentence representations in set $S$, thus achieving aligned language representations ($A$).
Algorithm 1 Language Representation and Alignment Process
0:  Dataset $D$, URIEL database $U$, Language Model $\Theta$, mBERT representations $M$
0:  Aligned Language Representations $A$
  $F_{U}\leftarrow$ ExtractFeatures($U$)
  $S\leftarrow\{\}$
  for each sentence $x$ in $D$ do
     $H_{x}\leftarrow$ GetLastHiddenStates($x$, $\Theta$)
     $N_{x}\leftarrow$ CountAttentionMasked($x$)
     $R_{x}\leftarrow\frac{\textsc{Sum}(H_{x})}{N_{x}}$
     $S\leftarrow S\cup\{R_{x}\}$
  end for
  $W,b\leftarrow$ InitializeParameters()
  for each training epoch do
     $P_{U}\leftarrow(W\times S)+b$
     $loss\leftarrow$ ComputeLoss($P_{U}$, $F_{U}$)
     $W,b\leftarrow$ UpdateParametersWithConstraint($W,b,loss$)
  end for
  $A\leftarrow\{\}$
  for each sentence representation $s$ in $S$ do
     $A_{m}\leftarrow(W\times s)+b$
     $A\leftarrow A\cup\{A_{m}\}$
  end for
Note that there may be discrepancies between the scales of the standard classification loss and the URIEL loss. To address this, we introduce an optional hyperparameter, denoted as $\lambda$, to scale the URIEL loss appropriately. However, finding this scaling factor requires another hyperparameter search. Therefore, we propose a new method for dynamically balancing the classification and URIEL losses using two dynamic scaling approaches.
Dynamic Scaling Approaches
In addition to the fixed scaling factor, we also explore dynamic adjustment of this scaling factor at each training step. This aims to maintain a balance between the classification and URIEL losses, and even considers making the scale trainable. The final loss formula when training with LinguAlchemy is given by:
$\mathcal{L}=\lambda_{cls}*\mathcal{L}_{cls}+\lambda_{uriel}*\mathcal{L}_{uriel}(Z,U).$ (2)
We define two methods to implement dynamic scaling:
  1. 1.
    AlchemyScale: This method dynamically adjusts the scaling factor $\lambda$ during training. It is initiated with scaling factors set relative to the mean of initial losses. Furthermore, these factors are updated periodically using an Exponential Moving Average (EMA) method to balance between different loss components.
  2. 2.
    AlchemyTune: Here, $\lambda$ is conceptualized as a trainable parameter within the model’s architecture. Initialized as part of the model’s parameters and optimized during the training process. This method applies the scaling factors to loss components, then an additional mini_loss is computed to represent the deviation of the sum of scaling factors.
Both methods aim to enhance model performance by dynamically and intelligently scaling loss components, with the first method relying on predefined, periodically updated scaling mechanisms, and the second integrating the scaling factor into the model’s learning parameters for adaptive adjustments.
4 Experiment Setting
Datasets
In our experiments, we use MASSIVE Dataset FitzGerald et al. (2023), which is a comprehensive collection of multilingual data incorporating intent classification tasks. We split MASSIVE into 25 languages that are “seen" during finetuning and the rest 27 languages that are “unseen", which we exclusively used for evaluation. This splitting is based on the language adapters availability as outlined in the prior research of Pfeiffer et al. (2020a), which we utilized in the AdapterFusion experiment for our baseline model. For a detailed breakdown of the languages used, including their respective families, genera, and script can be found in Appendix A.
Additionally, we incorporate the MasakhaNews Dataset Adelani et al. (2023), consisting of news article classification across several African languages. This dataset tests our models against diverse journalistic styles and complex syntactic structures. For our experiments, the training languages are amh, eng, fra, hau, swa, orm, and som, while the testing languages include ibo, lin, lug, pcm, run, sna, tir, xho, and yor. Lastly, we also utilize the SemRel2024 Dataset Ousidhoum et al. (2024) for semantic relatedness task in low-resource languages. We use this dataset to evaluate LinguAlchemy’s semantic understanding and relationship extraction capabilities. We train using the languages amh, arq, ary, eng, esp, hau, kin, mar, and tel. The test set includes afr, amh, arb, arq, ary, eng, esp, hau, hin, ind, kin, and pan.
Models
Our study employs two widely used and resource-efficient multilingual language models: Multilingual BERT Base (mBERTBase) and XLM-RoBERTa Base (XLM-RBase). In our training process, we use a learning rate of $5\times 10^{-5}$, train for 30 epochs, and measure performance based on accuracy for MASSIVE and MasakhaNews, and Pearson correlation for SemRel. Each training takes at most 5 hours using a single A100 GPU.
Model Language Category MASSIVE MasakhaNews SemRel
Base Ours Base Ours Base Ours
mBERT Low 48.43 66.93 ($\uparrow 18.50$) 46.25 72.70 ($\uparrow 26.45$) -0.26 -0.14 ($\uparrow 0.12$)
Medium 55.96 64.42 ($\uparrow 8.47$) 42.94 64.38 ($\uparrow 21.45$) 0.16 0.21 ($\uparrow 0.05$)
High 79.66 66.81 ($\downarrow 12.85$) 77.89 73.74 ($\downarrow 4.16$) 0.02 -0.03 ($\downarrow 0.05$)
XLM-R Low 78.15 80.90 ($\uparrow 2.76$) 47.86 79.31 ($\uparrow 31.46$) 0.08 0.40 ($\uparrow 0.32$)
Medium 80.31 80.15 ($\downarrow 0.16$) 55.87 75.79 ($\uparrow 19.92$) 0.49 0.37 ($\downarrow 0.12$)
High 86.82 80.86 ($\downarrow 5.96$) 68.14 73.79 ($\uparrow 5.65$) 0.26 0.41 ($\uparrow 0.15$)
Table 2: Performance comparison of mBERT and XLM-R models across different language categories and benchmarks.
5 Results and Discussion
5.1 LinguAlchemy Performance
Our results as shown in Table 5 reveal that LinguAlchemy excels across all languages in the MasakhaNews dataset, including those not encountered during the pretraining of mBERT (*) and XLM-R ($\dagger$). LinguAlchemy further demonstrates notable improvements on the Semantic Relatedness dataset, showing its ability to adapt to languages with distinct typological characteristics from the training corpus. In this experiment, we opted not to compare our method against the baseline used in the Semantic Relatedness paper because LaBSE is not zero-shot; it was pretrained with sentence similarity tasks contrasting our method’s conditions. Moreover, we excluded the MAD-X experiment from the MasakhaNews evaluation because MAD-X’s parameter-efficient approach differs fundamentally from our full finetuning approach. Collectively, these insights suggest that our LinguAlchemy can generalize across varied linguistic attributes.
Additionally, we applied the same procedure to MASSIVE dataset and the results are summarized in Table 6. We compared our method with zero-shot generalization, where the model is fully tuned on seen languages and then tested on unseen languages (referred to as Full FT in the Table). Furthermore, we explored AdapterFusion Pfeiffer et al. (2021a) as another baseline. AdapterFusion has shown better adaptation to unseen languages than naive zero-shot generalization. Unfortunately, many language adapters that we need for AdapterFusion is not available for XLM-R.
From Table 6, it is shown that LinguAlchemy achieves better generalization for unseen languages. We observed a significant improvement for mBERT and a modest average improvement for the stronger XLM-R model. For mBERT, LinguAlchemy can significantly increase performance in truly unseen languages of am-ET, km-KH, mn-MN, in which mBERT has never seen during the pre-training stage nor fine-tuning. These findings show that LinguAlchemy can be useful in truly zero-shot settings. While LinguAlchemy significantly boosts performance in weaker languages such as cy-GB or sw-KE, it can occasionally degrade results in languages with already strong zero-shot performance, particularly evident in XLM-R where it tends to flatten results to the 80-82% range.
Dataset Language UNK %
MASSIVE am-ET 6.79
km-KH 3.81
vi-VN 0.35
Other languages <0.1
SemRel amh 3.43
hau 0.60
ary 0.44
Other languages <0.4
MasakhaNews pcm 0.43
eng 0.16
Other languages <0.1
Table 3: UNK percentages in different datasets, illustrating the prevalence of unknown tokens that LinguAlchemy successfully manages.
Despite the variations in performance, the potential of LinguAlchemy is particularly clear in scenarios where zero-shot performance is inherently weak. Our hypothesis is that the model indirectly leverages familiar scripts encountered during pretraining and helps its ability to effectively handle UNK tokens. Advances in models using byte-level tokenization units theoretically reduce or eliminate OOV tokens; however, our evaluations across the MASSIVE, MasakhaNews, and SemRel2024 datasets, as shown in Table 3, confirm that UNK tokens have a minimal impact, thus showing the robustness of LinguAlchemy in such environments. For contexts where UNK token rates are high, the solution might be orthogonal to our approach, requiring further improvement in the base models or tokenizers that could later be integrated with LinguAlchemy.
5.2 Effect of Scaling URIEL Loss
The classification and URIEL losses operate on different scales. Simply adding these losses together would cause the model to prioritize the loss with the larger magnitude. During the early stages of training, we observe that the classification loss is approximately ten times larger than the URIEL loss. In this section, we explore the impact of various scaling factors applied to the URIEL loss.
Figure 3: Average performance of unseen languages under various URIEL loss scaling factors.
Constant Scaling
We investigate the effect of consistently scaling the URIEL loss by different factors. The results are illustrated in Figure 3. Notably, because we utilize the scale-invariant optimizer AdamW, there is no risk of gradients becoming excessively large due to high loss values.
As observed in Figure 3, for mBERT, the performance on unseen languages improves significantly when scaling the URIEL loss by 10x, achieving the highest performance of 68.68%. However, as the scaling factor increases further (e.g., 25x, 50x, and 100x), the performance starts to decline, indicating that overly emphasizing the URIEL loss has diminishing returns or even negative effects. Without any scaling (0x), the model performs poorly, showing the importance of scaling the URIEL loss.
For XLM-R, the trend is more stable. The performance fluctuates slightly across different scaling factors but remains generally consistent, with the highest performance achieved at the 10x scaling factor (80.32%). Larger scaling factors (e.g., 50x, 100x) do not lead to substantial improvements and may even cause minor drops in performance. This suggests that while scaling helps, XLM-R is less sensitive to the URIEL loss scaling than mBERT.
Overall, a scaling factor of 10$\times$ appears to give the best balance between classification and URIEL losses.
Dynamic and Trainable Scaling
Introducing a scaling factor adds another tunable hyperparameter, which can complicate the training process. Ideally, we seek a balanced weighting between the classification and URIEL losses. Instead of exhaustively testing various scaling factors, an adaptive scaling approach is more cost-effective and advantageous. Here, we explore two strategies: dynamic scaling and trainable scaling factors. The results of these approaches are presented in Table 4.
URIEL Scaling mBERT XLM-R
Constant 10$\times$ 64.68 80.32
AlchemyScale 62.97 80.43
AlchemyTune 63.24 79.10
Table 4: Performance comparison across different URIEL scaling methods.
Interestingly, these dynamic scaling methods do not significantly outperform a constant scaling factor. Specifically, a 10$\times$ scaling achieves the best performance for mBERT, while dynamic scaling only marginally outperforms the 10$\times$ scaling for XLM-R. Therefore, in scenarios with limited computational resources, a 10$\times$ scaling factor is recommended. However, with more computational capacity, exploring different scaling factors may yield marginal gains.
[…]
Table 6: Performance of LinguAlchemy in MASSIVE dataset for unseen languages. For languages in *, mBERT has never seen the languages during pre-training.
5.3 Generalization Across Language Family
We investigate LinguAlchemy across language families to further analyze the generalization capabilities of BERT and XLM-R models. We perform our experiment by splitting the languages in MASSIVE according to their language families and train the model on a subset of language families while testing on the rest, unseen language families. We explore on including different subset of language families, as seen in the Appendix (Table 7).
As illustrated in Figure 4, LinguAlchemy demonstrates generalization towards these unseen language families. Perhaps unsurprisingly, adding more subset of diverse languages improves generalization performance. Notably, the inclusion of the Afro-Asiatic language group—consisting of languages such as “am-ET", “ar-SA", and “he-IL", each featuring unique scripts—has significantly enhanced performance from the second to the third training group iteration. This improvement underscores LinguAlchemy’s capability to adapt to scripts not presented during the initial training or fine-tuning phases, illustrating its robustness in generalizing across different scripts.
The performance of both models, combined with LinguAlchemy underscores the advantage of including a broader spectrum of languages within training groups for enhanced model generalization. However, the impact of this diversity is not uniform across all language families: While some consistently benefit from the expansion of training data, others do not, and shows that merely increasing the volume of data from the same family may not necessarily improve performance. This inconsistency highlights the potential limitations within the models’ capacity to learn and generalize the linguistic features specific to certain language families. Consequently, our observation shows that the degree of generalization varies among different language families. This suggests that while some may significantly profit from these models’ capabilities, others may require more tailored strategies to gain similar performance improvement.
5.4 Seen Language Performance
While LinguAlchemy consistently improves performance across unseen languages, we note some inconsistencies concerning the performance of seen languages. In MASSIVE, we observe a noticeable performance drop in seen languages, while in contrast, we still see a massive gain in MasakhaNews and the performance of SemRel seems to be unaffected. The compiled results can be seen in Table 4.
Figure 4: Model performance across language families. Dotted lines indicates language families used in training in some of the training stages (solid dots for active use–refer to Table 7), and solid grey lines for families unseen in all training stages, with variance shown in shading.
In 4, we compare the performance of the BERT and XLMR models across different language families. The dotted lines represent the language families that were included in training (i.e., seen languages), and the solid gray lines represent language families that were unseen in training. Notably, variance for unseen languages is indicated by the shaded areas. This figure shows that while both models generally improve across training groups, the performance of certain language families, particularly seen languages, varies. For instance, in the case of the BERT model, languages from the Sino-Tibetan family demonstrate relatively poor performance across training stages compared to other families, even though they are part of the seen group. Meanwhile, XLMR shows a more consistent performance boost across language families, including unseen ones, though Indo-European languages perform better overall.
As MasakhaNews focuses on extremely low-resource languages, we hypothesize that despite being exposed during pretraining, the models’ performance remains low even with standard fine-tuning methods. Hence, LinguAlchemycan significantly improve performance, even by  18% in the low-resource languages. For high-resource languages, traditional fine-tuning is a better choice. We are investigating why LinguAlchemydoes not help with some languages and how to enhance the performance of some seen languages as part of our future work. Nevertheless, our method still proves beneficial in under-resourced settings where multilingual models typically perform poorly.
6 Conclusion
We introduced LinguAlchemy, a novel approach that demonstrates strong performance across 30+ unseen languages on intent classification and semantic relatedness tasks. Our method hinges on the integration of linguistic knowledge through the URIEL vectors, enhancing the language model’s ability to generalize across a diverse set of languages. We also proposed AlchemyScale and AlchemyTune, which employs a hyperparameter search for the URIEL scaling factor. This is achieved by two key strategies: (1) weight-averaging classification and URIEL loss, and (2) learning to balance the scale between classification and URIEL loss. LinguAlchemy achieves a massive performance improvement on low-resource languages in multiple downstream tasks including intent classification ($\uparrow 18.50$), news classification ($\uparrow 31.46$), and semantic relatedness ($\uparrow 0.32$).
Limitations
LinguAlchemy enhances performance across many unseen languages in intent classification, yet it faces limitations. Performance on seen languages is less than ideal, indicating room for improvement through methods like weight freezing. Also, better generalization appears to reduce accuracy in seen languages, pointing to a need for balanced approaches. Currently, the research is limited to intent classification, and expanding to other NLP tasks could reveal more about its versatility. Moreover, the choice of URIEL features—syntax, geography, language family—is theoretically sound, as discussed in Section 3, but empirical tests with different features might refine the model further. Overcoming these limitations could greatly improve the generalizability and effectiveness of multilingual NLP models.
Appendix A Languages in Dataset
The MASSIVE Dataset, also known as the Multilingual Amazon SLU Resource Package (SLUPR), offers a comprehensive collection of approximately one million annotated utterances for various natural language understanding tasks such as slot-filling, intent detection, and Virtual Assistant performance evaluation. It is an extensive dataset that includes 51 languages, 60 intents, 55 slot types, and spans 18 different domains. The dataset is further enriched with a substantial amount of English seed data, comprising 587k training utterances, 104k development utterances, and 152k test utterances.
Train Group Lang. Family Languages Num. Languages
1 Indo-European af-ZA, bn-BD, ca-ES, cy-GB, da-DK, de-DE, el-GR, en-US, es-ES, fa-IR, fr-FR, hi-IN, hy-AM, is-IS, it-IT, lv-LV, nb-NO, nl-NL, pl-PL, pt-PT, ro-RO, ru-RU, sl-SL, sq-AL, sv-SE, ur-PK 26
2 Dravidian Train Group 1 + kn-IN, ml-IN, ta-IN, te-IN 30
3 Afro-Asiatic Train Group 2 + am-ET, ar-SA, he-IL 33
4 Sino-Tibetan Train Group 3 + my-MM, zh-CN, zh-TW 36
Unseen Languages sw-KE, km-KH, vi-VN, id-ID, jv-ID, ms-MY, tl-PH, ja-JP, ka-GE, ko-KR, mn-MN, th-TH, az-AZ, tr-TR, fi-FI, hu-HU 16
Table 7: Language family distribution used in the language family generalization experiment (§5.3). The "others unseen" category includes additional language families not incorporated in the training set that we use as an “unseen" testbed.
Code Name Script Genus Code Name Script Genus
ar-SA Arabic Arab Semitic is-IS Icelandic Latn Germanic
bn-BD Bengali Beng Indic ka-GE Georgian Geor Kartvelian
el-GR Greek Grek Greek km-KH Khmer Khmr Khmer
en-US English Latn Germanic lv-LV Latvian Latn Baltic
es-ES Spanish Latn Romance ml-IN Malayalam Mlym Southern Dravidian
fa-IR Persian Arab Iranian nb-NO Norwegian Latn Germanic
fr-FR French Latn Romance ro-RO Romanian Latn Romance
he-IL Hebrew Hebr Semitic sl-SI Slovenian Latn Slavic
hu-HU Hungarian Latn Ugric ur-PK Urdu Arab Indic
hy-AM Armenian Armn Armenian zh-CN Mandarin Hans Chinese
id-ID Indonesian Latn Malayo-Sumbawan zh-TW Mandarin Hant Chinese
Table 8: Statistics and description of the dataset used Xu et al. (2022). The dataset used is a subset of the MASSIVE dataset, selecting 25 different seen languages.
Code Name Script Genus Code Name Script Genus
af-ZA Afrikaans Latn Germanic my-MM Burmese Mymr Burmese-Lolo
am-ET Amharic Ethi Semitic nl-NL Dutch Latn Germanic
az-AZ Azerbaijani Latn Turkic pl-PL Polish Latn Slavic
cy-GB Welsh Latn Celtic pt-PT Portuguese Latn Romance
da-DK Danish Latn Germanic ru-RU Russian Cyrl Slavic
de-DE German Latn Germanic sq-AL Albanian Latn Albanian
fi-FI Finnish Latn Finnic sv-SE Swedish Latn Germanic
hi-IN Hindi Deva Indic sw-KE Swahili Latn Bantoid
ja-JP Japanese Jpan Japanese ta-IN Tamil Taml Southern Dravidian
kn-IN Kannada Knda Southern Dravidian te-IN Telugu Telu South-Central Dravidian
ko-KR Korean Kore Korean th-TH Thai Thai Kam-Tai
mn-MN Mongolian Cyrl Mongolic vi-VN Vietnamese Latn Viet-Muong
ms-MY Malay Latn Malayo-Sumbawan
Table 9: Statistics and description of the dataset used Xu et al. (2022). The dataset used is a subset of the MASSIVE dataset, selecting 27 different unseen languages.
Appendix B Language Family Experiment
Tables 10 and 11 provide a comprehensive analysis of language family performance across different training groups. These tables compare the accuracy percentages of the Multilingual BERT and XLM-RoBERTa models, respectively. The results displayed in the tables elucidate the models’ capabilities in generalizing from the training data to unseen languages. A clear trend that can be observed is the improvement in performance as the training groups progress from 1 to 4, which suggests that the models benefit from exposure to a wider variety of language families during training. The ’Average’ row at the bottom of each table indicates the mean accuracy across all language families, providing an insight into the overall performance enhancement achieved by each model with incremental training diversity.
Language Family Train Group 1 Train Group 2 Train Group 3 Train Group 4
Afro-Asiatic 52.82% 52.93% 61.26% 61.00%
Atlantic-Congo 65.71% 68.08% 70.62% 71.79%
Austroasiatic 64.77% 66.78% 69.72% 70.16%
Austronesian 66.88% 68.66% 72.06% 72.19%
Dravidian 64.74% 67.97% 70.93% 71.41%
Indo-European 67.50% 68.61% 72.53% 72.95%
Japonic 72.11% 71.98% 75.80% 75.67%
Kartvelian 68.91% 68.89% 72.46% 72.32%
Koreanic 64.80% 66.46% 70.04% 69.91%
Mongolic-Khitan 63.11% 66.44% 69.71% 69.59%
Sino-Tibetan 62.65% 66.29% 68.79% 70.33%
Tai-Kadai 63.52% 67.89% 70.23% 71.34%
Turkic 54.69% 56.91% 63.54% 64.05%
Uralic 71.49% 71.27% 75.33% 75.15%
Average 65.54% 67.07% 71.04% 71.43%
Table 10: Multilingual BERT Performance of Language Families Across Training Groups
Language Family Train Group 1 Train Group 2 Train Group 3 Train Group 4
Afro-Asiatic 75.74% 76.23% 85.56% 85.39%
Atlantic-Congo 70.86% 72.38% 83.24% 82.73%
Austroasiatic 74.85% 76.04% 83.91% 83.59%
Austronesian 78.94% 79.83% 84.77% 84.69%
Dravidian 81.49% 82.20% 85.41% 85.43%
Indo-European 80.31% 81.21% 83.26% 83.47%
Japonic 80.21% 81.36% 82.67% 83.15%
Kartvelian 80.40% 81.53% 82.79% 83.27%
Koreanic 79.74% 80.91% 82.14% 82.61%
Mongolic-Khitan 79.54% 81.00% 82.20% 82.65%
Sino-Tibetan 79.25% 81.00% 82.14% 82.58%
Tai-Kadai 79.08% 80.81% 81.90% 82.35%
Turkic 79.20% 80.90% 81.96% 82.39%
Uralic 79.24% 80.91% 81.92% 82.47%
Average 79.45% 80.48% 83.44% 83.62%
Table 11: XLM-RoBERTa Performance of Language Families Across Training Groups
Appendix C Appendix: Language Identification (LID) Experiments
This section presents the results of comprehensive language identification experiments performed across a variety of popular language detection models. The evaluation is detailed in two distinct tables:
Table 12 displays the performance of traditional language identification models such as LID-Fasttext, CLD3, CLD2, langid, and LangDetect across multiple languages within the MASSIVE dataset. These results illustrate the effectiveness of each model in correctly identifying the language of given text samples.
Table 13 focuses on the accuracy of multilingual language models, specifically XLM-R and mBERT, alongside adaptations using the MAD-X framework with embeddings from FastText and CLD3. This evaluation aims to show how these advanced models perform in the task of language identification, especially in comparison to more specialized LID tools.
Language LID-Fasttext CLD3 CLD2 langid LangDetect
ar-SA 94.25 86.45 81.58 91.78 94.13
bn-BD 99.72 97.52 89.57 96.93 99.76
de-DE 97.70 88.59 89.73 92.83 82.54
el-GR 99.68 96.91 99.77 99.84 99.64
en-US 98.61 79.44 93.43 93.96 87.82
es-ES 96.20 78.24 73.14 86.87 86.55
fi-FI 97.70 92.91 92.90 92.08 96.09
fr-FR 98.35 87.53 85.23 94.77 94.80
hi-IN 98.44 88.21 97.83 87.94 93.54
hu-HU 98.54 92.24 93.89 95.34 96.71
hy-AM 99.90 98.37 99.92 99.17 0.00
id-ID 87.20 65.86 73.54 72.68 89.32
is-IS 89.93 92.64 90.88 92.97 0.00
ja-JP 99.41 96.63 99.04 99.11 96.23
jv-ID 24.75 68.10 0.00 22.04 0.00
ka-GE 99.56 98.49 99.95 99.65 0.00
ko-KR 99.50 98.47 99.03 99.96 99.36
lv-LV 90.73 90.06 95.25 94.33 97.32
my-MM 99.93 96.90 99.97 0.00 0.00
pt-PT 92.17 83.42 77.39 77.74 84.05
ru-RU 99.27 84.48 82.35 83.79 91.32
vi-VN 98.41 95.85 97.26 98.62 99.53
zh-CN 97.55 98.07 84.33 99.64 0.00
zh-TW 95.76 94.19 0.03 99.31 0.00
Average 93.89 89.57 83.17 86.31 66.20
Table 12: Per language results of language identification evaluation in MASSIVE.
Language XLMR mBERT MAD-X
MAD-X
w/ FastText
MAD-X
w/ CLD3
ar-SA 79.32 78.35 75.72 71.92 67.79
bn-BD 83.25 80.23 78.61 76.36 74.95
de-DE 85.54 83.59 81.81 79.49 76.90
el-GR 85.07 81.74 80.93 79.56 78.51
en-US 88.16 86.45 85.78 83.89 83.15
es-ES 86.18 84.97 82.58 80.97 76.43
fi-FI 85.24 82.55 82.55 79.86 77.07
fr-FR 86.48 86.11 83.69 82.35 80.03
hi-IN 84.63 82.38 80.73 78.14 72.73
hu-HU 85.68 82.65 81.57 80.13 76.40
hy-AM 84.23 81.20 80.43 78.78 77.91
id-ID 86.52 84.67 82.01 76.03 69.30
is-IS 84.16 82.21 80.40 71.49 73.57
ja-JP 85.78 84.70 83.22 82.04 81.27
jv-ID 81.20 81.57 78.58 45.70 59.68
ka-GE 79.19 75.25 73.23 70.85 70.17
ko-KR 85.51 84.30 82.99 81.14 80.56
lv-LV 84.73 82.18 82.08 74.58 74.95
my-MM 82.18 78.01 78.48 76.36 74.98
pt-PT 86.35 85.27 83.59 80.56 77.77
ru-RU 86.65 83.96 83.52 81.74 75.45
vi-VN 86.48 83.32 82.52 79.72 78.61
zh-CN 85.41 85.24 84.23 53.09 52.69
zh-TW 83.73 82.55 81.27 52.79 52.45
Average 84.65 82.64 81.27 74.90 73.47
Table 13: Per language accuracy score of multilingual language models in MASSIVE.